# Reputation in the Shadow of Exit

### v5

@@ 6 Discussion
Second, suppose instead that the inequality fails for some $\alpha\in(0,1)$ but that it holds for $\alpha=1$. This implies first that it is possible to sustain the Markov equilibrium where the long-run player plays $c^{*}$ at all histories. Moreover, there must exist some $\overline{\alpha}\in(0,1)$ with $a\in\mathcal{N}(\alpha c^{*})$ and
@@
by the intermediate value theorem, since the incentive compatibility constraint is continuous in $\alpha$. This implies that there exists a Markov equilibrium of the following form:
@@
    If $\mu(h^{t})(\omega_{R})\neq\{\mu_{0}(\omega_{R}),1\}$, then have $\sigma_{L}(h^{t})=c^{*}$ and $\{\sigma_{i}(h^{t})\}\in\mathcal{N}(c^{*})$ be chosen to give the long-run player $V_{*}$.
@@
    If $\mu(h^{t})(\omega_{R})$, play static Nash $\sigma(h^{t})$ giving the long-run player $\overline{V}_{0}$.
@@
    If $\mu(h^{t})(\omega_{R})=\mu_{0}(\omega_{R})$, have $a=\{\sigma_{i}\}\in\mathcal{N}(\alpha c*)$ chosen so that
@@
    $\mu_{0}(\omega_{R})\sigma(h^{t})(c^{*})+(1-\mu_{0}(\omega_{R}))=\alpha$
    Noting this is well-defined so long as $\mu_{0}(\omega_{R})>1-\alpha$.
It is clear that at all histories, the short-run player is best replying, as is the long-run player at all histories outside of $\varnothing$. Moreover, at the first history $h^{t}=\varnothing$, by choice of $\alpha$, the long-run player is indifferent between $c^{*}$ and 0 and thus is willing to randomize given $a$. Hence this is a Markov perfect equilibrium. This gives the long-run player an equilibrium payoff of
@@
where the first inequality follows from monotonicity of the game and the second follows from the fact the incentive compatibility condition holds at $\alpha=1$ where $a_{*}=a^{*}(c^{*})$ is the largest element of $\mathcal{N}(c^{*})$, since again by the monotonicity assumption this is the choice which minimizes the stage-game payoff. This implies the result.
Finally, suppose the inequality fails when $\alpha=1$. Then there cannot be an equilibrium where the long-run player imitates $c^{*}$ for sure. The only other pure Markov equilibrium gives the long-run player their minimax payoff and implies the result. If instead the long-run player mixes in a Markov equilibrium at some on-path history, they secure (for some $a\in\mathcal{N}(\alpha c^{*})$ and some $\alpha\in(0,1)$) a payoff of at most
@@
because dependence on $a$ is suppressed in the likelihood ratio and $|\mathcal{N}(0)|$ is a singleton so $\underline{V}_{0}=\overline{V}_{0}$. These are the same incentive compatibility constraints. Hence the two statements are equivalent.
@@
Fix $\varepsilon>0$ and preferences $(u,v)$. Fix some action $c$ and $a^{*}(c)$ be some best reply to it where $v(a^{*}(c),c)>0$, which exists by the nontriviality assumption. Set $f(a^{*}(c),c)=1-\eta,f(a^{*}(c),0)=\eta$. Let $\underline{a}=a^{*}(0)\in\mathcal{N}(0)$ be such that $(\underline{a},0)$ gives the long-run player payoff $\underline{V}_{0}$. Set moreover $f^{*}(\underline{a},0)=f^{*}(\underline{a},c)=\eta$. Note for $\eta$ sufficiently small, playing $c$ must be preferred to playing 0 since
$\frac{v(a^{*}(c),c)(1-\eta)}{1-(1-\eta)}\geq\frac{v(a,0)\eta}{1-\eta}$
for any $a\in\mathcal{N}(0)$. This in particular implies for all $\eta>0$, there is an optimal commitment posture $c^{*}\neq 0$ and in particular one which guarantees a payoff of at least $v(a^{*}(c),c)>0$ in each period conditional on surviving.
$f(a^{*}(c^{*}(\eta)),c^{*})[v(a^{*}(c^{*}(\eta)),c^{*})+V_{*}(\delta,\eta)]\geq f(a^{*},0)[v(a^{*}(0)+\underline{V}_{0}]$
The term inside the brackets on the right hand side is independent of $\eta$, and the left hand side by assumption is bounded from below by
$f(a^{*}(c),c)[v(a^{*}(c),c)+V_{*}(\eta)]$
as $\eta$ vanishes. But then our condition is implied by the requirement that
$\frac{f(a^{*}(c),c)}{f(a^{*},0)}[v(a^{*}(c),c)+V_{*}(\eta)]\geq[v(a^{*}(0)+\delta V_{*}]$
$\iff$ $\frac{1-\eta}{\eta}[v(a^{*}(c),c)+V_{*}(\eta)]\geq[v(a^{*}(0)+\underline{V}_{0}]$
which is clearly true as $\eta\to 0$. Hence under this survival probability function $f^{*}$, $c^{*}$ optimality is satisfied for all $\eta$ sufficiently small.
  1. (1)
    Have the long-run player play 0 in the first period.
  2. (2)
    Have the short-run player’s play to $a^{*}(\mu_{0}(\omega_{c^{*}})c^{*}))$, i.e. best respond to their prior belief supposing the rational type plays 0.
  3. (3)
    Regardless of the long-run player’s action, given them a payoff of $V^{*}$ tomorrow.
Clearly $V^{*}$ is sustainable by the folk theorem tomorrow; and because the rational long-run player’s action today does not affect their future continuation payoff, they will play their dominant action 0. Finally, Condition (2) is exactly the condition for short-run players to myopically best reply to this strategy by the long-run player. As $\mu_{0}(\omega_{R})\to 1$, short-run players (by upper hemi-continuity) converge to an action in $\mathcal{N}(0)$. In particular, this implies $\underline{a}$ must be short-run players a (weakly) lower payoff than $\lim\limits_{\alpha\to 0}v(a^{*}(\alpha c^{*}),c^{*})$.
This equilibrium strategy profile gives the long-run player a payoff of exactly
which diverges as $\varsigma\to\infty$. ∎
PROOF OF PROPOSITION 4
Proof.
The first equivalence. Rank $F\succ A$ and $In\succ Out$. Then $f(a,c)$ and $v(a,c)$ are strictly decreasing for each fixed $c$, and $\mathcal{N}(\alpha\text{Fight})$ is decreasing in $\alpha$, so the game is monotone. Thus the first statement follows from Theorem 1 once we verify $c^{*}$-incentive compatibility. First, note that
@@
Clearly the expression on the right hand side is decreasing in $\varepsilon$ and satisfied as $\varepsilon\to 0$ for fixed $\varphi$, and so in particular the intermediate value theorem furnishes the necessary $\varepsilon$ (possibly $\varepsilon=1$) for the result to hold.
The second part of the result follows by constructing the type of equilibrium in Proposition 3, taking $\varepsilon\to 0$ and noting that the long-run player will not fight in the first period because the “exit probability” is sunk. ∎
@@
We have that
$\frac{\partial U_{i}}{\partial c}=F_{\theta}(F_{x}(x|x)|x)-1\text{ and }\frac{\partial U_{i}}{\partial x}=(1+c)\left(F_{\theta}(F_{x}(x|x)|x)\cdot f_{x}(x|x)\cdot\frac{\partial}{\partial 2}F_{\theta}(F_{x}(x|x)|x)\cdot\frac{\partial}{\partial 2}F_{x}(x|x)\right)$
@@
$\lim\limits_{\varsigma\to 0}F_{x}\left(\frac{1}{1+c^{*}}\bigg|x\right)\to\begin{cases}0\text{ if }x<\frac{1}{1+c^{*}}\\ 1\text{ if }x>\frac{1}{1+c^{*}}\end{cases}$
@@ B.3. The Main Result
Say the long-run player asymptotically satisfies (weak) $c^{*}$ incentive compatibility if there exists $\overline{\delta}<1$ such that for all $\delta>\overline{\delta}$ and each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,

### v6

@@ 6 Discussion
Second, suppose instead that the inequality fails for some $\beta\in(0,1)$ but that it holds for $\beta=1$. This implies first that it is possible to sustain the Markov equilibrium where the long-run player plays $c^{*}$ at all histories. Moreover, there must exist some $\overline{\beta}\in(0,1)$ with $a\in\mathcal{N}(\beta c^{*})$ and
@@
by the intermediate value theorem, since the incentive compatibility constraint is continuous in $\beta$. This implies that there exists a Markov equilibrium of the following form:
@@
    If $\mu(h^{t})(\omega_{R})\not\in\{\mu_{0}(\omega_{R}),1\}$, then have $\sigma_{L}(h^{t})=c^{*}$ and $\{\sigma_{i}(h^{t})\}\in\mathcal{N}(c^{*})$ be chosen to give the long-run player $V_{*}$.
@@
    If $\mu(h^{t})(\omega_{R})=1$, play static Nash $\sigma(h^{t})$ giving the long-run player $\overline{V}_{0}$.
@@
    If $\mu(h^{t})(\omega_{R})=\mu_{0}(\omega_{R})$, have $a=\{\sigma_{i}\}\in\mathcal{N}(\beta c*)$ chosen so that
@@
    $\mu_{0}(\omega_{R})\sigma(h^{t})(c^{*})+(1-\mu_{0}(\omega_{R}))=\beta$
    Noting this is well-defined so long as $\mu_{0}(\omega_{R})>1-\beta$.
It is clear that at all histories, the short-run player is best replying, as is the long-run player at all histories outside of $\varnothing$. Moreover, at the first history $h^{t}=\varnothing$, by choice of $\beta$, the long-run player is indifferent between $c^{*}$ and 0 and thus is willing to randomize given $a$. Hence this is a Markov perfect equilibrium. This gives the long-run player an equilibrium payoff of
@@
where the first inequality follows from monotonicity of the game and the second follows from the fact the incentive compatibility condition holds at $\beta=1$ where $a_{*}=a^{*}(c^{*})$ is the largest element of $\mathcal{N}(c^{*})$, since again by the monotonicity assumption this is the choice which minimizes the stage-game payoff. This implies the result.
Finally, suppose the inequality fails when $\beta=1$. Then there cannot be an equilibrium where the long-run player imitates $c^{*}$ for sure. The only other pure Markov equilibrium gives the long-run player their minimax payoff and implies the result. If instead the long-run player mixes in a Markov equilibrium at some on-path history, they secure (for some $a\in\mathcal{N}(\beta c^{*})$ and some $\beta\in(0,1)$) a payoff of at most
@@
because dependence on $a$ is suppressed in the likelihood ratio and $|\mathcal{N}(0)|$ is a singleton so $\underline{V}_{0}=\overline{V}_{0}$. These are the same incentive compatibility constraints modulo an equality. Hence the two statements are generically equivalent.
@@
The first equivalence. Rank $F\succ A$ and $In\succ Out$. Then $f(a,c)$ and $v(a,c)$ are strictly decreasing for each fixed $c$, and $\mathcal{N}(\beta\text{Fight})$ is decreasing in $\beta$, so the game is monotone. Thus the first statement follows from Theorem 1 once we verify $c^{*}$-incentive compatibility. First, note that
@@
Clearly the expression on the right hand side is decreasing in $\varepsilon$ and satisfied as $\varepsilon\to 0$ for fixed $\varphi$, and so in particular the intermediate value theorem furnishes the necessary $\varepsilon$ for the result to hold.
The second part of the result follows by constructing the following equilibrium:
  1. (1)
    The long-run player accommodates in the first period. Short-run players myopically best reply by entering.
  2. (2)
    In all future periods, players give the long-run player a payoff of $V_{*}$.
The long-run player chooses Accomodate in the first period as $\varphi,\varepsilon$ vanish, and the long-run player’s payoff is bounded from above by $2\varepsilon$, which vanishes. ∎
@@
Note $F_{x}(x|x)=\mathbb{P}(x_{j}<x|x)$ which in the symmetry noise model is $\frac{1}{2}$. We thus have that We have that
$\frac{\partial U_{i}}{\partial c}=F_{\theta}\left(\frac{1}{2}|x\right)-1\text{ and }\frac{\partial U_{i}}{\partial x}=(1+c)\left(\frac{\partial}{\partial 2}F_{\theta}\left(\frac{1}{2}|x\right)\right)$
@@
$\lim\limits_{\varsigma\to 0}F_{x}\left(\frac{1}{1+c^{*}}\bigg|x\right)\to\begin{cases}1\text{ if }x<\frac{1}{1+c^{*}}\\ 0\text{ if }x>\frac{1}{1+c^{*}}\end{cases}$
@@ B.3. The Main Result
Say the long-run player asymptotically satisfies (weak) $c^{*}$ incentive compatibility if there exists $\overline{\delta}<1$ such that for all $\delta>\overline{\delta}$ and each $a\in\mathcal{N}(\beta c)$ for $\beta\in[0,1]$,